OpenStudy (anonymous):

How do I compute the following calculation with the different units of energy? 70kg x 9.8m/s^2 x 2m

OpenStudy (anonymous):

actually u'll get the unit of energy after the multipliction

OpenStudy (anonymous):

kg m/s^2 is the unit of force and 'm' is the unit of displacement

OpenStudy (anonymous):

so answer will be (98*14) joule

OpenStudy (anonymous):

how did you get 14?

OpenStudy (anonymous):

70*2* 9.8 = 140* 9.8 = 14 * 98 = 1372

OpenStudy (anonymous):

so, just for my general knowledge as this is my first physics class in college... I can just multiply out the numbers, and then change the symbol accordingly depending on what it is that i'm multiplying? I took Bus Calc last semester so i'm used to not being able to combine terms if they aren't the same lol

OpenStudy (anonymous):

in this expression 1st term is mas 2nd term is acceleration 3rd term is displacement so mass times acceleration is force and frce times displacement is work done...which has the unit of energy

OpenStudy (anonymous):

ahhh gotcha.. thank you so much! :)

OpenStudy (anonymous):

:)